OpenStudy (anonymous):

What is the solution to the system of equations 5x-3y=-9 2x-5y=4

OpenStudy (jackky):

I also need help on that too can any one help me please?

OpenStudy (anonymous):

and me :P

OpenStudy (anonymous):

ohh and me!! xp

OpenStudy (anonymous):

Do you know substitution or elimination?

OpenStudy (anonymous):

yes partially

OpenStudy (anonymous):

Which do you want to use?

OpenStudy (anonymous):

umm im not sure :P

OpenStudy (anonymous):

5x-3y=-9 2x-5y=4 2(5x-3y)=-9*2 10x - 6y = -18 -5(2x-5y)=4*-5 -10x + 25y = -20 10x - 6y = -18 -10x + 25y = -20 ---------------------------- 0 + 19y= -38 19y/19 = -38/19 y = -2 Now we know y = -2. Take -2 and put it into one of the other equations and solve for x. 5x-3(-2)=-9 Can you finish and find x?

OpenStudy (anonymous):

Elimination Take note of the coefficients of \(x\)\[ \color{red}5x-3y=-9\\ \color{blue}2x-5y=4 \]Multiply the equations so that they are the same: \[ \color{blue}2\cdot(\color{red}5x-3y=-9)\\ \color{red}5 \cdot (\color{blue}2x-5y=4) \]Then it becomes: \[ 10x-6y=-18\\ 10x-25y=20 \]Now you can subtract the second equation from the first: \[ (-6-(-25))y = -18 -20\\ 19y=-38 \]Now you can solve for \(y\)\[ y=-2 \]Then put that into the equations above: \[ 5x-3(-2)=-9 \]And you can solve for \(x\).

OpenStudy (anonymous):

sorry it wouldnt load... but would I do the same thing for the first one?

OpenStudy (anonymous):

for the second one I mean

OpenStudy (anonymous):

What second one?

OpenStudy (anonymous):

for x

OpenStudy (anonymous):

.... :/

OpenStudy (anonymous):

Well, you have to isolate \(x\).

OpenStudy (anonymous):

\[ 5x+6=-9 \]Move the \(6\) over: \[ 5x=-15 \]So... what times \(5\) is \(-15\)?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

so (-2, 3)

OpenStudy (anonymous):

@Jackky @lopez09

OpenStudy (jackky):

idk any of this srry im startin to learn it

OpenStudy (anonymous):

(-2, 3)

OpenStudy (anonymous):

or at least thats what I think it is but @wio hasn't said anything yet

OpenStudy (anonymous):

You can always plug it into both equations to see if it is true

OpenStudy (anonymous):

However, \[ 5(3) = 15\neq -15 \]

OpenStudy (anonymous):

\[ 5x=-15\\ x = \frac{-15}{5}=-3 \]

OpenStudy (anonymous):

so negative 3 not 3

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you